# FloVD: Optical Flow Meets Video Diffusion Model for Enhanced Camera-Controlled Video Synthesis

### v1

@@ 5.3 Comparison
We first compare the camera controllability of our method against MotionCtrl [33] and CameraCtrl [11]. Both MotionCtrl and CameraCtrl were trained on RealEstate10K [43], which provides no object motions but a wider range of camera motions than our full dataset. For a comprehensive comparison, we evaluate four versions of our model. Specifically, we train FVSM on either our internal dataset or RealEstate10K, but without utilizing the ground-truth camera parameters available in RealEstate10K. We also include variants of our model with and without OMSM as RealEstate10K contains only static scenes without moving objects. In this evaluation, OMSM is trained using our internal dataset.
@@
Our method also achieves better performances for the benchmark datasets of object motion synthesis quality (Pexels-small, Pexels-medium, and Pexels-large), as reported in Tab. 2. While CameraCtrl exhibits significantly degraded quality for large object motions (Pexels-large), our method achieves substantially better results for all three benchmark datasets. MotionCtrl often fails to follow input camera parameters, synthesizing videos where the viewpoint remains close to the input image. This may lead to good FVD scores, as the synthesized videos align well with the minimal camera movement present in most benchmark videos. However, as shown in Fig. 6, MotionCtrl often produces visual artifacts in the synthesized videos. More visual examples for these artifacts can be found in the additional qualitative comparison of the supplemental document. Furthermore, our method, which employs a timestep sampling strategy from the EDM framework [16], outperforms previous methods across all metrics, as demonstrated in Tab. 3 and Tab. S1 of the supplemental document.
@@
5.4 Further Analysis
Training Timestep Pexels-random
Data Strategy FVD ($\downarrow$) FID ($\downarrow$) IS ($\uparrow$)
Baseline RE10K QTS 157.99 39.61 11.19
+ OMSM RE10K QTS 104.92 36.33 11.51
+ large-scale data Internal QTS 91.55 35.66 11.63
Baseline RE10K EDM 148.42 39.92 11.23
+ OMSM RE10K EDM 95.31 36.90 11.43
+ large-scale data Internal EDM 80.74 35.65 11.73
@@ 5.4 Further Analysis
Our method generates camera flow maps using the 3D structure estimated from an input image, then feeds them to FVSM. To better understand our framework, Fig. 8 presents visualizations of warped images using the estimated 3D structure and camera parameters, alongside their associated video synthesis results produced by FVSM. As shown in the figure, 3D-based image warping may introduce distortions and holes, yet still provides realistic-looking images. This result indicates that leveraging the 3D structure can serve as a powerful hint for camera-controllable video synthesis. Fig. 8 also shows that our flow-conditioned video synthesis successfully produces realistic-looking results that closely resemble the warping results, but without artifacts such as distortions and holes.
Timestep sampling strategy.
As stated in Sec. 5.1, we adopt the quadratic timestep sampling strategy (QTS) for better camera controllability, instead of the timestep sampling strategy of the EDM framework [16] used in Stable Video Diffusion [3]. Our model using QTS shows slightly compromised video synthesis quality compared to our model using EDM (Tab. 3). Nevertheless, our model with QTS still demonstrates better performance than the previous methods (Tab. 2). Moreover, QTS enhances the camera controllability of FVSM (Tab. 4). This improved camera controllability results from increased chances of training with highly noised data, allowing our model to effectively leverage flow conditions for structural content generation.
mRotErr () mTransErr mCamMC
Ours w/ EDM 1.70 0.0983 0.1010
Ours w/ QTS 1.43 0.0869 0.0887
Table 4: Timestep sampling strategies for camera controllability.
@@ 5.5 Applications
Our framework using FVSM enables temporally-consistent video editing at no extra cost. A video can be edited as follows. First, we obtain optical flow maps from the video by using an off-the-shelf optical flow estimator [28]. We then edit the first frame of the input video using an off-the-shelf image editing tool, e.g., InfEdit [36]. We synthesize a video by using FVSM with the edited first frame and estimated optical flow maps as inputs. As shown in Fig. 9, this simple strategy using our framework can produce temporally-consistent video editing results, thanks to our optical-flow-based approach.
@@
Thanks to its 3D-awareness, our framework enables complex camera controls such as the dolly zoom. The dolly zoom is a camera technique where the camera moves forward while simultaneously adjusting the zoom in the opposite direction. To achieve this, we synthesize videos with control over both the camera’s intrinsic and extrinsic parameters. Fig. 10 shows a synthesized video with the dolly zoom effect, where the subject remains a similar size while the background appears to converge inward. Notably, our framework achieves this without being trained on video datasets containing camera intrinsic parameters that vary across frames.
@@
Figure 10: Video synthesis results with the dolly zoom effect.
@@ 6 Conclusion
This paper proposed FloVD, a novel optical-flow-based video diffusion model for camera controllable video generation. Since existing methods require a training dataset with ground-truth camera parameters, they are mainly trained on the restricted datasets that primarily consist of static scenes, leading to video synthesis with unnatural object motion. Unlike previous methods, our method leverages optical flow maps to represent both camera and object motions, enabling the use of arbitrary training videos without ground-truth camera parameters. Moreover, our method facilitates detailed camera control by leveraging background motions of optical flow, which encodes 3D correlation across different viewpoints. Our extensive experiments demonstrate that FloVD provides realistic video synthesis with natural object motion and accurate camera control.
@@
Our method is not free from limitations. Errors from both the object motion synthesis model and the semantic segmentation model may result in unnatural object motion in the synthesized videos. The estimation error of segmentation model can be alleviated through user interaction by providing point prompts for object regions. Our future work will involve a seamless integration of camera and object motions to synthesize more natural videos.

### v2

@@ 5.3 Comparison
We first compare the camera controllability of our method against MotionCtrl [33] and CameraCtrl [9]. Both MotionCtrl and CameraCtrl were trained on RealEstate10K [43], which provides no object motions but a wider range of camera motions than our full dataset. For a comprehensive comparison, we evaluate four versions of our model. Specifically, we train FVSM on either our internal dataset or RealEstate10K, but without utilizing the ground-truth camera parameters available in RealEstate10K. We also include variants of our model with and without OMSM, as RealEstate10K contains only static scenes without moving objects. In this evaluation, OMSM is trained using our internal dataset.
@@
Our method also achieves better performances for the benchmark datasets of object motion synthesis quality (Pexels-small, Pexels-medium, and Pexels-large), as reported in Tab. 2. While CameraCtrl exhibits significantly degraded quality for large object motions (Pexels-large), our method achieves substantially better results for all three benchmark datasets. MotionCtrl often fails to follow input camera parameters, synthesizing videos where the viewpoint remains close to the input image. This may lead to good FVD scores, as the synthesized videos align well with the minimal camera movement present in most benchmark videos. However, as shown in Fig. 6, MotionCtrl often produces visual artifacts in the synthesized videos. These artifacts are also evidenced by the degraded FID scores for MotionCtrl. More visual examples of the visual artifacts can be found in the supplemental document. In addition, by employing a timestep sampling strategy from the EDM framework [14], our method outperforms previous methods across all metrics (Tab. S1 in the supplemental document).
@@
5.4 Analysis
In the following, we provide an ablation study of our main components and a detailed mechanism of FVSM. Refer to Sec. 5 in the supplemental document for further analysis.
@@ 5.4 Analysis
Training Pexels-random
Data FVD ($\downarrow$) FID ($\downarrow$) IS ($\uparrow$)
Baseline RE10K 157.99 39.61 11.19
+ OMSM RE10K 104.92 36.33 11.51
+ large-scale data Internal 91.55 35.66 11.63
@@
Our method generates camera flow maps using the 3D structure estimated from an input image and feeds them to FVSM. To better understand our framework, Fig. 8 presents visualizations of warped images using the estimated 3D structure and camera parameters, alongside their associated video synthesis results produced by FVSM. As shown in the figure, 3D-based image warping may introduce distortions and holes, yet still provides realistic-looking images. This result indicates that leveraging the 3D structure can serve as a powerful hint for camera-controllable video synthesis. Fig. 8 also shows that our flow-conditioned video synthesis successfully produces realistic-looking results that closely resemble the warping results, but without artifacts such as distortions and holes.
@@ 5.5 Applications
Our framework using FVSM enables temporally-consistent video editing at no extra cost. Specifically, we first obtain optical flow maps from the input video and edit the first frame of the video. We then synthesize a video by using FVSM with the edited first frame and flow maps as inputs, producing temporally-consistent video editing results (Fig. 9).
@@
Thanks to its 3D awareness, our framework supports advanced camera controls such as the dolly zoom, which moves the camera forward or backward while simultaneously adjusting the zoom in the opposite direction. Fig. 1(Left) shows a synthesized video with the dolly zoom effect, where the subject remains a similar size while the background appears to converge inward. Notably, our framework accomplishes this without requiring training on video datasets with varying camera intrinsic parameters across frames.
@@ 6 Conclusion
This paper proposed FloVD, a novel optical-flow-based video diffusion model for camera-controllable video generation. Since existing methods require a training dataset with ground-truth camera parameters, they are mainly trained on restricted datasets that primarily consist of static scenes, leading to video synthesis with unnatural object motion. Unlike previous methods, our method leverages optical flow maps to represent both camera and object motions, enabling the use of arbitrary training videos without ground-truth camera parameters. Moreover, our method facilitates detailed camera control by leveraging background motions of optical flow, which encodes 3D correlation across different viewpoints. Our extensive experiments demonstrate that FloVD provides realistic video synthesis with natural object motion and accurate camera control.
@@
Our method is not free from limitations. Errors from both the object motion synthesis model and the semantic segmentation model may result in unnatural object motion in the synthesized videos. The estimation error of the segmentation model can be alleviated through user interaction by providing point prompts for object regions. Our future work will involve a seamless integration of camera and object motions to synthesize more natural videos.
Acknowledgment.
This work was supported by the Korea government (MSIT), through the IITP grant (Global Research Support Program in the Digital Field program, RS-2024-00436680; Development of VFX creation and combination using generative AI, RS-2024-00395401; Artificial Intelligence Graduate School Program (POSTECH), RS-2019-II191906) and NRF grant (RS-2023-00211658; RS-2024-00438532), and Microsoft Research Asia.